# Fermi-LAT Discovery of a Gamma-ray Outburst from the Peculiar Compact Steep Spectrum Radiogalaxy 3C 216

### v1

@@ 4 Discussion and conclusions
The simultaneous flaring activity observed in optical, UV, X-ray and gamma-rays between 2023-05-01 and 2023-05-09 provides a confirmation of the identification of the gamma-ray source with 3C 216. The TS map of 4FGL J0910.0+4257 in the total period is shown in Fig. 7. We marked in green the position of 3C 216, as reported by Petrov et al. (2005). Our results also confirm that the SSC process can well explain the production of gamma-ray outbursts from this type of radio sources. In VLBA archive [^6] https://science.nrao.edu/facilities/vlba/data-archive, there are data to study the morphology of the jet close to the central engine of the source. The results of our investigation could serve as foundation for future radio Target of Opportunity proposals if further high-energy flaring activity occurs.
@@
The Fermi LAT Collaboration acknowledges generous ongoing support from a number of agencies and institutes that have supported both the development and the operation of the LAT as well as scientific data analysis. These include the National Aeronautics and Space Administration and the Department of Energy in the United States, the Commissariat à l’Energie Atomique and the Centre National de la Recherche Scientifique / Institut National de Physique Nucléaire et de Physique des Particules in France, the Agenzia Spaziale Italiana and the Istituto Nazionale di Fisica Nucleare in Italy, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), High Energy Accelerator Research Organization (KEK) and Japan Aerospace Exploration Agency (JAXA) in Japan, and the K. A. Wallenberg Foundation, the Swedish Research Council and the Swedish National Space Board in Sweden. INFN and ASI personnel performed in part under ASI-INFN Agreements No. 2021-43-HH.0 The NASA Swift gamma-ray burst explorer is a MIDEX Gamma Ray Burst mission led by NASA with participation of Italy and the UK. This research has made use of the Smithsonian/NASA’s ADS bibliographic database. This research has made use of the NASA/IPAC NED database (JPL CalTech and NASA, USA). This research has made use of the archives and services of the Space Science Data Center (SSDC), a facility of the Italian Space Agency (ASI Headquarter, Rome, Italy). This research has made use of the XRT Data Analysis Software (XRTDAS) developed under the responsibility of the SSDC. Supported by Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations). Facilities: Fermi Gamma-ray Space TelescopeLATSwiftXRTUVOT

### v2

@@ 5 Conclusions
The simultaneous flaring activity observed in optical, UV, X-ray and gamma-rays between 2023-05-01 and 2023-05-09 provides a confirmation of the identification of the $\gamma$-ray source 4FGL J0910.0+4257 with 3C 216. The TS map of 4FGL J0910.0+4257 in the total period is shown in Fig. 7. We marked in green the position of 3C 216, as reported by Petrov et al. (2005). Our results also suggest that the SSC process can well explain the production of $\gamma$-ray outbursts from this type of radio sources, thus being more universally applicable across different AGN classes than previously thought. In VLBA archive [^7] https://science.nrao.edu/facilities/vlba/data-archive, there are data to study the morphology of the jet close to the central engine of the source. The results of our investigation could serve as foundation for future radio Target of Opportunity proposals if further high-energy flaring activity occurs.
Due to its overall steep radio spectrum and relatively compact morphology, 3C 216 is classified as a CSS source. In the youth scenario proposed by Fanti et al. (1995), these sources owe their compactness to a young age and represent an early evolutionary stage towards a fully developed radio galaxy size. CSS sources, and their even more compact (and younger) siblings (GHz Peaked Spectrum, or GPS, sources, O’Dea et al. 1991), were predicted to be good candidates for GeV emission, as they combine on sub-galactic scales both the presence of recently injected relativistic particles and abundant photon fields from the central regions of their hosts (Stawarz et al., 2008). However, only a handful of CSS/GPS sources have been detected individually (Migliori et al., 2016; Principe et al., 2020) and even a stacking analysis has not revealed a collective signal from this population (Principe et al., 2021).
Except for a few outstanding sources detected because of their extreme proximity, a significant contribution from Doppler beamed components, such as relativistic jets, is required for CSS and GPS sources to be detected in $\gamma$ rays. This is suggested by elements such as the position in the luminosity-photon index diagram, the identification with quasar hosts, and the presence of variability. While Principe et al. (2021) already pointed out all these characteristics for 3C 216, the flare analyzed here reveals a level of flux variability significantly more extreme than anything observed before in this source.
Acknowledgments
@@
The Fermi LAT Collaboration acknowledges generous ongoing support from a number of agencies and institutes that have supported both the development and the operation of the LAT as well as scientific data analysis. These include the National Aeronautics and Space Administration and the Department of Energy in the United States, the Commissariat à l’Energie Atomique and the Centre National de la Recherche Scientifique / Institut National de Physique Nucléaire et de Physique des Particules in France, the Agenzia Spaziale Italiana and the Istituto Nazionale di Fisica Nucleare in Italy, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), High Energy Accelerator Research Organization (KEK) and Japan Aerospace Exploration Agency (JAXA) in Japan, and the K. A. Wallenberg Foundation, the Swedish Research Council and the Swedish National Space Board in Sweden.
INFN and ASI personnel performed in part under ASI-INFN Agreements No. 2021-43-HH.0.
The NASA Swift gamma-ray burst explorer is a MIDEX Gamma Ray Burst mission led by NASA with participation of Italy and the UK. This research has made use of the Smithsonian/NASA’s ADS bibliographic database. This research has made use of the NASA/IPAC NED database (JPL CalTech and NASA, USA). This research has made use of the archives and services of the Space Science Data Center (SSDC), a facility of the Italian Space Agency (ASI Headquarter, Rome, Italy). This research has made use of the XRT Data Analysis Software (XRTDAS) developed under the responsibility of the SSDC.
Supported by Italian Research Center on High Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2 within the activities of Spoke 3 (Astrophysics and Cosmos Observations).
We also gratefully acknowledge Giacomo Principe, Filippo D’Ammando, Melissa Pesce-Rollins, Deirdre Horan, and anonymus external referees whose comments and suggestions contributed to shaping this work.
Facilities: Fermi Gamma-ray Space TelescopeLATSwiftXRTUVOT